OpenStudy (anonymous):

arthquake intensity is given by I = Io(10)m, where Io is the reference intensity and m is magnitude. A particular major earthquake of magnitude 6.0 is 90 times as intense as a particular minor earthquake. The magnitude, to the nearest tenth, of the minor earthquake is

OpenStudy (paxpolaris):

\[\Large I = I_0\cdot10^m\] if \(I_M\) is the major earthquakes intensity and \(I_x\) is is the minor earthquakes intensity \[\Large I_M= I_0\cdot10^6\]\[\Large I_x= I_0\cdot10^x\] and: \[\Large I_M= 90 \cdot I_x\]

OpenStudy (paxpolaris):

earthquakes =earthquake's

OpenStudy (paxpolaris):

\[\Large \implies{I_M \over I_x}=90\] \[\Large \implies{\cancel{I_0}\cdot10^6 \over \cancel{I_0}\cdot10^x}=90\]

OpenStudy (paxpolaris):

then take the log of both sides:\[\Large \implies 6-x=\log \left( 90 \right)\]\[\Large \implies \LARGE x=6-\log \left( 90 \right)\]